# ReliableMath: Benchmark of Reliable Mathematical Reasoning on Large Language Models

### v1

@@ 4.2 Dataset Analysis
In experiments, we observe that LLMs are prone to fail on removal questions or those with difficulty level $d$=1 on unsolvable problems. Therefore, we make some analysis of our constructed unsolvable data regarding different rewriting types and difficulty levels as in Fig. 5.
Rewriting Types (Removal & Contradiction)
In Sec. 3.2 and Fig. 2, we present two unsolvable rewriting types of removal and contradiction. As shown in Fig. 5 (a), we present the precision of unsolvable problems employing two rewriting methods (removal & contradiction) on several LLMs using reliable prompts. Generally, Prec.($\mathcal{U}$) of contradiction problems is larger than removal problems, indicating that LLMs are more adept at identifying the unsolvability of contradiction cases. As we observed in generated outputs, LLMs tend to fabricate missing conditions for removal problems, resulting in verbose yet hallucinated responses, thus yielding fewer successful responses.
@@
As described in Step 3 of Sec. 3.3, human experts think that problems with a difficulty level of 0 can be intuitively judged as unsolvable, whereas problems with a difficulty level of 1 require step-by-step reasoning to determine their unsolvability. In Fig. 5 (b), we showcase the precision of unsolvable problems with two difficulty levels (0: simple & 1: hard). Prec.($\mathcal{U}$) of unsolvable problems with difficulty level 1 is lower than problems of difficulty level 0 across all LLMs, suggesting that LLMs are consistently correlated with human experts in identifying unsolvability of problems. As in Appendix G, we have specified detailed principles for difficulty level evaluation to minimize variations arising from the subjectivity of human experts and ensure data quality.
@@ 5 Reliability Improvements
With the preceding analysis, since such relatively small LLMs exhibit worse reliability, we further propose a Supervised Fine-tuning-based (SFT) alignment strategy in Sec. 5.1 to improve LLM reliability, and conduct experiments to validate the effectiveness of the strategy in Sec. 5.2.
@@ 5.1 Alignment Strategy
Figure 6: Illustrations of (a) reliability alignment training data generation for both solvable and unsolvable problems, and (b) SFT-based alignment using the obtained training data.
To enhance the reliability of small LLMs $\mathcal{M}_{S}$, we first obtain alignment training data. Given an training set ${\mathcal{Q}_{a}}^{(t)}$ with open-source solvable mathematical problems, we synthesize unsolvable problems in ${\mathcal{Q}_{u}}^{(t)}$ using the unsolvable data construction process in Sec. 3.3  [^2] Due to the expensive cost of human checks on substantial training data, Step 3 is omitted. As demonstrated in Table 10, the pass rate of Step 3 human check is approximately 80.14%; therefore, the quality of synthesized unsolvable problems after Step 2 is regarded as applicable enough.. We then generate successful and refused responses for problems in ${\mathcal{Q}_{a}}^{(t)}$ and ${\mathcal{Q}_{u}}^{(t)}$.
Due to the limited ability of $\mathcal{M}_{S}$ to generate successful responses including high-quality reasoning paths of solvable problems or identify unsolvability of solvable problems, we employ an advanced teacher LLM $\mathcal{M}_{T}$ (DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability, )) for distillation (xu2024survey, ). As in Fig. 6 (a), we employ the standard prompt $p_{s}$ to distill from $\mathcal{M}_{T}$ to sample successful responses $\mathcal{S}$ for problems in ${\mathcal{Q}_{a}}^{(t)}$ and the reliable prompt $p_{r}$ for ${\mathcal{Q}_{u}}^{(t)}$, with a sampling number $K=4$ to ensure that most problems yield at least one $\mathcal{S}$ response. To prevent LLM from refusing to answer questions it can solve, for ${\mathcal{Q}_{a}}^{(t)}$, we first infer on small LLMs to select questions that are incorrectly answered and add to unknown problem set ${\mathcal{Q}_{a}}^{(n)}$. Then we employ rejection sampling (yuan2023scalingrelationshiplearningmathematical, ) on $\mathcal{M}_{S}$ to sample model-specific refusal responses $\mathcal{R}$ for ${\mathcal{Q}_{u}}^{(n)}$ and ${\mathcal{Q}_{n}}^{(n)}$ with $K=1$ and a 2-shot refusal prompt $p_{f}$ to accelerate sampling. Since $\mathcal{S}$ is more preferable than $\mathcal{F}$, we only need a few refusal responses to enable $\mathcal{M}_{S}$ to appropriately learn the refusal capability. The generated data are incorporated into the training set ${\mathcal{D}}^{(t)}$ to align with the reliability for $\mathcal{M}_{s}$ using SFT as in Fig. 6 (b). Prompt templates are in Appendix I.1.
5.2 Alignment Experiments
Table 4: Results of in-domain (AIME, AMC, MATH in average) and out-of-domain (Minerva) Precision and Prudence on both solvable and unsolvable problems on two small LLMs (Distill-1.5B, Qwen2.5-1.5B) using standard and reliable prompts, and our proposed alignment strategy, where DeepSeek-R1 is the teacher model. $\dagger$ denotes that significant improvements are obtained using the alignment method over both reliable and standard prompting methods.
@@ 5.2 Alignment Experiments
Prec.($\mathcal{A}$) Prud.($\mathcal{A}$) Prec.($\mathcal{U}$) Prud.($\mathcal{U}$) Prec.($\mathcal{A}$) Prud.($\mathcal{A}$) Prec.($\mathcal{U}$) Prud.($\mathcal{U}$)
@@
standard 0.540 0.000 0.000 0.000 0.260 0.000 0.000 0.000
reliable 0.516 0.000 0.003 0.000 0.220 0.000 0.039 0.000
alignment 0.663 0.086 0.231 0.104 0.300 0.100 0.190 0.078
@@
standard 0.531 0.000 0.000 0.000 0.180 0.000 0.000 0.000
reliable 0.507 0.000 0.000 0.000 0.160 0.000 0.000 0.000
alignment 0.684 0.094 0.272 0.114 0.320 0.072 0.184 0.087
[…]
@@
We test the in-domain (ID) performance on AIME, AMC, and MATH as well as the out-of-domain (OOD) performance on Minerva of the ReliableMath dataset. Results in Table 4 suggest that our alignment strategy can effectively improve the reliability of two LLMs on both ID and OOD tasks, with significant improvements of $0.2\sim 0.4$ in Prec.($\mathcal{U}$) over two baseline prompting methods. Surprisingly, advancements in Prud.($\mathcal{A}$) and Prud.($\mathcal{U}$) indicate that aligned LLMs acquire the refusal ability on both solvable and unsolvable problems, which is rare when using prompting methods.
@@ 6 Conclusion
This work introduces a ReliableMath benchmark to systematically investigate LLM reliability on reasoning tasks. The benchmark formulates the reliability evaluation on mathematical reasoning tasks, as well as develops a dataset comprising open-source solvable math problems and high-quality unsolvable problems synthesized from our proposed data construction workflow including human evaluation. Experiments conducted on various reasoning and instruction LLMs derive several key findings and analyses. Finally, we propose an alignment strategy which can significantly enhance the reliability of small LLMs. We believe this study can serve as a strong benchmark to inspire more valuable research to develop more reliable LLMs for future work.
Appendix A Ethics Statement
Appendix B Limitations
The investigation of the ReliableMath benchmark is solely limited to mathematical reasoning, while other reasoning tasks, like code, also require exploration of reliability. This constraint primarily arises from the reliability definition may vary in different reasoning tasks. A comprehensive benchmark of LLM reliability applicable to all reasoning tasks is challenging. Future research endeavors will extend the ReliableMath benchmark to a wider range of benchmarks on different reasoning tasks.
In addition, excessive optimization of LLMs’ reliability may impair their creativity and decrease performance in some open-ended generation tasks. To provide users with more reliable responses or avoid misleading them, LLMs may prematurely terminate reasoning even for problems where they could have provided partially useful responses.
Appendix C Potential Social Impact
Appendix D Definition of Notations
@@
$\boldsymbol{y}_{i}$ The ground-truth answer to $\boldsymbol{x}_{i}$ in $\mathcal{D}_{a}$.
$\boldsymbol{\hat{y}}_{i}$ The LLM generated response to $\boldsymbol{x}_{i}$.
@@
${\boldsymbol{\hat{x}}_{i}^{j}}$ The $j$-th rewritten question of $\boldsymbol{x}_{i}$.
${\boldsymbol{\hat{c}}_{i}^{j}}$ The rewritten condition by model verification of ${\boldsymbol{\hat{x}}_{i}^{j}}$ where ${\boldsymbol{\hat{c}}_{i}^{j}}$ entails ${\boldsymbol{c}_{i}^{j}}$.
${\boldsymbol{\hat{a}}_{i}^{j}}$ The unsolvable reason analysis by model verification of ${\boldsymbol{\hat{x}}_{i}^{j}}$.
${{d}_{i}^{j}}$ The difficulty level by human evaluation of ${\boldsymbol{\hat{x}}_{i}^{j}}$.
@@
Appendix E Related Works
E.1 LLM Reliability
The primary function of large language models is to generate reliable responses for users, meaning that the content should be factually accurate, informative, consistent, and trustworthy to users [28, 25, 13]. Reliability is a foundational requirement for LLM alignment because unreliable outputs would negatively undermine almost all LLM applications and mislead users [48], such as decision-making [63] and problem-solving [14] contexts.
Although the definition and categorization of LLM reliability are many-sided and depend on specific tasks, a unified perspective is that LLMs should be able to identify questions that fall outside their knowledge scope or are inherently meaningless and unanswerable [57, 3, 46]. The most direct approach to enhancing the reliability of knowledge-based tasks is to teach LLMs to refuse to answer questions that are beyond their knowledge boundaries [62], more generally, express what they know and do not know, rather than providing hallucinated answers with confidence. Recent works tackle the problem by aligning LLMs with factuality and uncertainty [52, 55, 50, 27, 64, 53].
However, current research on reliability primarily focuses on knowledge-intensive tasks [57, 3], while studies on reasoning tasks remain scarce. Such methods or evaluations are model-specific. For more complex reasoning tasks, users not only expect the LLMs to refuse unknown questions, but also that LLMs can determine data-specific solvability. There is a void of reliable benchmarks and datasets, widely acknowledged evaluation standards, and effective improvement methods in this field.
E.2 Reasoning LLM
Recent advanced LLMs like OpenAI-o1 [38] and DeepSeek-R1 [10] have gained significant research interests in improving their reasoning capabilities [5, 36]. Such LLMs, also referred to reasoning LLMs, are designed to emulate the slow and deliberate Chain-of-Thought (CoT) [49] reasoning steps (System 2 thinking), while previous foundation and instruction-tuned LLMs like GPT-4o [35], DeepSeek-V3 [9], and Llama-3 [15] execute fast and heuristic-driven generation (System 1 thinking) [26, 18]. The improvements of reasoning LLMs have been primarily focused on the following three aspects: 1) Many studies have employed reinforcement learning techniques like PPO [40] and GRPO [10] to incentivize LLMs to explore more effective reasoning strategies [41, 8]; 2) Researchers also endeavor to develop carefully curated, high-quality reasoning datasets to promote reasoning performance [56, 34]. 3) Test-time scaling law [42] also inspires to elicit LLMs to produce long-CoT contents with various search algorithms during inference time [12].
Although reasoning LLMs have showcased remarkable performance on various benchmarks, they have revealed several shortcomings. Numerous investigations [47, 6, 39, 29] identified the “overthinking” issue where reasoning LLMs always generate massive and verbose thinking steps even for a simple or meaningless query, resulting in potential hallucinations and severe redundancy of tokens. Additionally, the negative impact of long-form reasoning on reliability remains underexplored and is still subject to critical skepticism.
E.3 Mathematical Tasks for LLMs
Mathematical tasks have long been regarded as effective proxies for both enhancing and assessing the reasoning capabilities of LLMs [20, 32, 43]. These tasks require LLMs to comprehend the semantics and symbols in the question, engage in problem-solving processes step by step, and present the final answers. Moreover, the verifiability of the final answer facilitates straightforward evaluation and reward design for rule-based RL methods [7, 41, 54].
In previous years, GSM8K [7] and MATH [16] have been utilized as prominent mathematical benchmarks for LLM training and evaluation. Recently, to further facilitate the reasoning capabilities of LLMs, researchers have paid attention to more challenging, Olympiad- or college-level mathematical problems like AIME [45] and CollegeMath [43]. Therefore, extensive mathematical training corpora have been proposed [23, 59, 4, 30].
Moreover, concerns are raised that LLMs genuinely master reasoning versus pattern memorization, and many studies have sought to evaluate true mathematical reasoning abilities through various verification methods [61, 33, 24, 11]. Nevertheless, prior work has not considered the scenario in which LLMs attempt to generate solutions when confronted with unknown or unsolvable problems, posing potential risks of hallucination and undermining reliability. There is also a lack of corresponding datasets containing unsolvable mathematical problems for thorough assessments.
Appendix F Dataset Details
F.1 Solvable Data Collection

### v2

@@ 4.3 Dataset Analysis
It is observed that LLMs are prone to fail on removal questions or those with difficulty level dd=1 of unsolvable cases, we thus make some analysis of our constructed unsolvable data regarding different rewriting schemes and difficulty levels as in Fig. 5.
Rewriting Schemes (Removal & Contradiction)
We present two unsolvable rewriting schemes of removal and contradiction in Sec. 3.2. As in Fig. 5 (a), we showcase the results of Succ.($\mathcal{U}$) by on two rewriting schemes on several LLMs using reliable prompts. Generally, Succ.($\mathcal{U}$) of contradiction problems is larger than removal problems, indicating that LLMs are more adept at identifying the unsolvability of contradiction cases. As we observed in generated outputs, LLMs tend to fabricate missing conditions for removal problems, resulting in verbose yet hallucinated responses, thus yielding fewer successful responses.
@@
As in Step 3 of Sec. 3.3, human experts label problems that are intuitively judged as unsolvable with a difficulty level of 0, whereas problems with a difficulty level of 1 require step-by-step reasoning to determine the unsolvability. In Fig. 5 (b), we demonstrate Succ.($\mathcal{U}$) with two difficulty levels (0: simple & 1: hard). Succ.($\mathcal{U}$) with difficulty level $d=1$ is lower than problems of $d=0$ across all LLMs, suggesting that LLMs are consistently correlated with human experts in identifying unsolvability of problems.
@@ 5 Reliability Improvements
With the preceding analysis, since such relatively small LLMs exhibit worse reliability, we further propose an alignment strategy in Sec. 5.1 to critically identify unsolvable problems, and conduct experiments to improve LLM reliability on both math and knowledge QA tasks in Sec. 5.3 and 5.2.
@@ 5.1 Alignment Strategy
Figure 6: Illustrations of (a) reliability alignment training data generation for both solvable and unsolvable problems, (b) SFT-based alignment using the obtained training data, and (c) DPO-based alignment on the obtained SFT model.
@@
Succ.($\mathcal{A}$) Refu.($\mathcal{A}$) Succ.($\mathcal{U}$) Refu.($\mathcal{U}$) Succ.($\mathcal{A}$) Refu.($\mathcal{A}$) Succ.($\mathcal{U}$) Refu.($\mathcal{U}$)
$s_{p}$ $s_{o}$ $s_{p}$ $s_{o}$ $s_{p}$ $s_{o}$
@@
standard 41.22 45.05 0.00 1.89 0.00 0.00 4.74 0.07 17.08 17.76 0.12
reliable 39.63 42.18 0.00 2.17 1.44 0.00 3.69 2.50 23.33 4.64 5.45
alignment-sft 48.24 48.89 9.58 17.15 22.69 9.56 3.16 10.87 31.36 25.23 12.48
alignment-dpo 46.00 46.64 5.11 21.87 27.22 8.31 1.57 8.37 25.44 30.26 7.88
@@
standard 40.26 41.86 0.00 2.74 0.00 0.00 0.13 0.00 11.70 0.06 0.00
reliable 38.98 39.62 0.00 2.18 0.00 0.00 0.26 0.07 11.88 0.12 0.00
alignment-sft 55.27 56.23 7.99 26.13 25.05 9.25 2.51 9.86 16.74 21.65 10.14
alignment-dpo 53.05 53.99 4.79 27.95 29.31 8.60 0.43 6.88 22.14 24.45 7.12
Table 5: Results of Success Rate (Succ.) and Refusal Rate (Refu.) on both in-domain math task and out-of-domain knowledge QA task including both solvable and unsolvable problems. Two relatively small LLMs of Distill-1.5B and Qwen2.5-1.5B are employed using standard and reliable prompts, as well as our proposed alignment strategy including both SFT and DPO. DeepSeek-R1 is the teacher model. $\dagger$ denotes that significant improvements are obtained using the alignment method over both reliable and standard baseline prompting methods.
To enhance the reliability of small LLMs $\mathcal{M}_{S}$, we first obtain alignment training data. Given the training set $\mathcal{Q}_{a}$} with open-source solvable math problems, we synthesize unsolvable problems in $\mathcal{Q}_{u}$} using the unsolvable data construction process in Sec. 3.3.333Due to the expensive cost of human checks on substantial training data, Step 3 is omitted. As demonstrated in Table 2, the pass rate of Step 3 human check is approximately 80.14%; therefore, the quality of synthesized unsolvable problems after Step 2 is regarded as applicable enough. We then generate successful and refused responses for problems in $\mathcal{Q}_{a}$} and $\mathcal{Q}_{u}$}.
Due to the limited ability of $\mathcal{M}_{S}$ to generate successful responses including high-quality reasoning paths of solvable problems or identify unsolvability of solvable problems, we employ an advanced teacher LLM $\mathcal{M}_{T}$ (DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability)) for distillation (xu2024survey). As in Fig. 6 (a), we employ the standard prompt $p_{s}$ for $\mathcal{Q}_{a}$} and the reliable prompt $p_{r}$ for $\mathcal{Q}_{u}$} to distill successful responses $\mathcal{S}$ from $\mathcal{M}_{T}$, with a sampling number $K=4$ to ensure that most problems yield at least one $\mathcal{S}$ response. To prevent LLM from refusing to answer questions it can solve, for $\mathcal{Q}_{a}$}, we first infer on small LLMs to select questions that are incorrectly answered and add to the unknown problem set $\mathcal{Q}_{n}$}. Then we employ rejection sampling (yuan2023scalingrelationshiplearningmathematical) on $\mathcal{M}_{S}$ to sample model-specific refusal responses $\mathcal{R}$ for $\mathcal{Q}_{n}$} and $\mathcal{Q}_{u}$} with $K=1$ and a 2-shot refusal prompt $p_{f}$ to accelerate sampling. Since $\mathcal{S}$ is more preferable than $\mathcal{F}$, we only need a few refusal responses to enable $\mathcal{M}_{S}$ to appropriately learn the refusal capability. The generated data are incorporated into the training set ${\mathcal{D}}^{(t)}$ to align with the reliability for $\mathcal{M}_{s}$ as follows. Prompt templates are in Supplement 15.
In Fig. 6 (b), the constructed reliable supervision data ${\mathcal{D}}^{(t)}$ is used for Supervised Fine-Tuning (SFT) to train the small LLM $\mathcal{M}_{s}$ to critically identify the unsolvability or refuse to answer while keeping their ability on solvable problems. Following the initial SFT phase, as in Fig. 6 (c), we sample and synthesize high-quality preference data based on the model’s outputs, which is subsequently utilized to further refine the model using the Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage) method, progressively aligning its behavior with the desired reliability preferences.
5.2 Alignment Setup
We conduct experiments on unreliable LLMs of Qwen2.5-1.5B and Distill-1.5B for reasoning reliability enhancements. We incorporate solvable problems from open-source AIME, AMC, and MATH training data into $\mathcal{Q}_{a}$}. To avoid potential data leakage due to the fundamental relation between train and test data which could be a confounder, questions in the $\mathcal{Q}_{a}$} with high similarity to the solvable test set $\mathcal{D}_{a}$ were filtered out as follows: For each $\boldsymbol{x}_{t}\in{\mathcal{Q}_{a}}$, we calculate the cosine similarity scores $\{s_{i}\}_{i=1}^{N}$ of $\boldsymbol{x}_{t}$ with all $\{\boldsymbol{x}_{i}|\boldsymbol{x}_{i}\in\mathcal{D}_{a}\}_{i=1}^{N}$, and if $\max\{s_{i}\}_{i=1}^{N}>0.7$, $\boldsymbol{x}_{t}$ is removed. The threshold $0.7$ is set empirically which filters out the top 10% problems with high similar scores.
When constructing unsolvable problems of $\mathcal{Q}_{u}$}, only one condition is extracted to maintain the size of $\mathcal{Q}_{u}$} comparable to $\mathcal{Q}_{a}$}. Distilled responses with reasoning steps are used to train the reasoning LLM Distill-1.5B, and we remove the reasoning steps to train the instruction LLM Qwen2.5-1.5B. The responses and problems are formatted and constituted to the alignment training set ${\mathcal{D}}^{(t)}$ ($|\mathcal{D}^{(t)}|=\text{10k}$). Statistics of training problems, the numbers of sampled $\mathcal{S}$ and $\mathcal{R}$ responses of $\mathcal{Q}_{a}$} and $\mathcal{Q}_{u}$}, and training details are all presented in Supplement 13.
LLMs after alignment are tested on both in-domain (ID) math tasks of ReliableMath dataset and out-of-domain (OOD) knowledge QA task derived from the KUQ dataset (amayuelas-etal-2024-knowledge). KUQ contains answerable knowledge QAs from open-source QA datasets like TriviaQA (joshi-etal-2017-triviaqa), as well as real-world unknown questions collected from crowd-source workers. Specifically, we omit categories of controversial and ambigious questions as they are not actually answerable but have various candidate answers. When inference, we also employ standard and reliable knowledge QA prompts for inference as in Supplement 15. Unlike math tasks, known knowledge questions typically do not involve explicit reasoning and can be directly evaluated using the F1 score on the outputs. Nonetheless, for unanswerable knowledge questions, it remains essential to assess whether LLMs’ reasoning process can identify the unsolvability aspects.
5.3 Alignment Experimental Results
@@
Reliability improvements suggest that our proposed alignment methods enable LLMs not only to simply identify the unsolvability of math problems, but also to critically think and identify their knowledge or capability boundary before answering any questions. As in Table 5, both of our alignment strategy using SFT and DPO can effectively improve the reliability of two LLMs on either ID mathematical or OOD real-world knowledge QA tasks, with significant improvements of $20\sim 30$ in Succ.($\mathcal{U}$) over two baseline prompting methods. Advancements in Refu.($\mathcal{A}$) and Refu.($\mathcal{U}$) indicate that aligned LLMs acquire the refusal ability on both solvable and unsolvable/unknown problems, which is rare when using prompting methods. In addition, although LLMs using DPO alignment exhibit larger gains on Succ.($\mathcal{U}$), they present inferior generalization on other reliability metrics compared with using SFT alignment.
6 Related Work
6.1 LLM Reliability
Reliability is a foundational requirement for LLM requiring generating factually correct, informative, and consistent responses to users (liu2023trustworthy; li2024surveyhonestylargelanguage; NEURIPS2024_0d99a8c0). A unified perspective of LLM reliability is that LLMs should be able to identify questions that fall outside their scopes or are unanswerable (yin-etal-2023-large; amayuelas-etal-2024-knowledge; wang-etal-2024-enhancing). Prior work enhances the reliability of knowledge tasks by aligning LLMs with their uncertainty to confidently answer known questions and refuse unknown questions (zhang-etal-2024-r) to mitigate hallucinations (xue2024ualignleveraginguncertaintyestimations; yang2024alignment; xu2024rejection; lin2024flamefactualityawarealignmentlarge; zheng2025enhancingllmreliabilityexplicit; xue2025mlingconfcomprehensivestudymultilingual). However, current research primarily focuses on knowledge tasks. There is a void of benchmark on LLM reliability on reasoning tasks.
6.2 Mathematical Reasoning on LLMs
Math tasks have long been regarded as effective proxies of reasoning capabilities of LLMs (koncel2016mawps; miao2021diverse; tang2024mathscalescalinginstructiontuning). These tasks require LLMs to comprehend the semantics and symbols in the question, engage in problem-solving processes step by step, and present the final answers. (shao2024deepseekmathpushinglimitsmathematical; yang2024qwen25mathtechnicalreportmathematical). Open benchmarks range from school-level GSM8K (cobbe2021training) and MATH (hendrycks2021measuring) to more challenging, Olympiad- or college-level AIME (aime_1983_2024) and CollegeMath (tang2024mathscalescalinginstructiontuning). Recent advanced LLMs like DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability), emulating the slow and deliberate deep Chain-of-Thought (CoT) (NEURIPS2022_9d560961) reasoning steps (System 2 thinking), have exhibited remarkable math reasoning abilities (chen2025reasoningerasurveylong) over previous foundation and instruction LLMs like DeepSeek-V3 (deepseekai2024deepseekv3technicalreport) which execute fast and heuristic-driven generation (System 1 thinking) (li202512surveyreasoning; loo2021system12). LLMs endowed with deep reasoning capabilities for solving complex math problems are commonly referred to as reasoning LLMs (chen2025reasoningerasurveylong). However, prior work has not considered the reasoning reliability evaluation of LLMs where unreliable LLMs fabricate solutions to unsolvable/unknown/unanswerable problems, thus leading to hallucination in reasoning steps.
@@ 7 Conclusion
This work introduces a ReliableMath benchmark to systematically investigate LLM reliability on reasoning tasks. The benchmark formulates the reliability evaluation on mathematical reasoning tasks, as well as develops a dataset comprising open-source solvable math problems and high-quality unsolvable problems synthesized from our proposed data construction workflow including expert check. Experiments conducted on a series of reasoning and instruction LLMs derive several key findings and analyses. Finally, we propose an alignment strategy which can enable LLMs to critically identify unsolvable cases and significantly enhance the reliability. We believe this study can serve as a strong benchmark to inspire more valuable research to develop more reliable LLMs for future work.
8 Definition of Notations
@@
$\boldsymbol{\hat{y}}_{i}$ The ground-truth answer to $\boldsymbol{x}_{i}$ in $\mathcal{D}_{a}$.
$\boldsymbol{y}_{i}$ The LLM generated response to $\boldsymbol{x}_{i}$.
@@
$\boldsymbol{\tilde{x}}_{i}^{j}$} The $j$-th rewritten question of $\boldsymbol{x}_{i}$.
$\boldsymbol{\tilde{c}}_{i}^{j}$} The rewritten condition by model verification of $\boldsymbol{\tilde{x}}_{i}^{j}$} where $\boldsymbol{\tilde{c}}_{i}^{j}$} entails $\boldsymbol{c}_{i}^{j}$}.
$\boldsymbol{\tilde{a}}_{i}^{j}$} The unsolvable reason analysis by model verification of $\boldsymbol{\tilde{x}}_{i}^{j}$}.
${d}_{i}^{j}$} The difficulty level by human evaluation of $\boldsymbol{\tilde{x}}_{i}^{j}$}.
@@
9 Dataset Details
9.1 Solvable Data Collection